OpenStudy (anonymous):

What is the square root of negative 100?

OpenStudy (anonymous):

10i

OpenStudy (anonymous):

Since you cannot sq.rt any negatives you must seperate it to this: \[\sqrt{-1} * \sqrt{100}\] since [\sqrt{-1\] is i and Sq.rt of 100 is 10

OpenStudy (anonymous):

10i

OpenStudy (anonymous):

So would the answer come out as a negative number or a positive number?

OpenStudy (anonymous):

\[\sqrt{-1}=i\] \[\sqrt{-100}=i \sqrt{10*10}=10i\]

OpenStudy (anonymous):

So what is the name of the symbol 100 goes inside when trying to come up with the square root?

OpenStudy (anonymous):

just to interrupt, i has no life of its own it is just another name for \[\sqrt{-1}\]so all you are saying is that \[\sqrt{-100}=\sqrt{100}\times \sqrt{-1}=10\sqrt{-1}\] and then replace \[\sqrt{-1}\] by \[i\]

OpenStudy (anonymous):

the negative sign is on the outside of the symbol the 100 is in. does that change the answer?

jimthompson5910 (jim_thompson5910):

\[\large -\sqrt{100}=-10\] \[\large \sqrt{-100}=10i\] \[\large -\sqrt{-100}=-10i\]

OpenStudy (anonymous):

okay.....so my answer is -10. thank you so much!!

jimthompson5910 (jim_thompson5910):

np, so I'm guessing there was only one negative in that problem then